# Inflection Point of Minimally Coupled Tachyonic Scalar Field

### v1

@@ Abstract
In this paper, we investigate the behaviour of a minimally coupled tachyonic scalar field at inflection points in an accelerating universe. We consider the different expansion factors and obtain potentials of tachyonic scalar field. Inflection points of homogeneous tachyonic scalar field is calculated for these potentials. We employ the tools of dynamical system analysis for the considered potentials and obtain the stable points.
@@ 1 Introduction
The cosmic acceleration is now an established phenomenon. As supported by several cosmological and astrophysical observations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] our universe is in the phase of accelerated expansion at the present stage and generally, it is believed that behind this accelerated expansion there is a mysterious component with negative pressure dominating $70\%$ of the universe. One of the good candidates that can explain this acceleration is ”Dark Energy” [11, 12, 13, 14, 15] which is introduced as modified matter in Einstein’s field equation which sits with the energy-momentum tensor. One of the remarkable features of dark energy is its negative pressure and repulsive gravity characteristics led by the equation of state $w_{de}<-1/3$. In alternative gravity theories, cosmic acceleration can be accounted for by modifications to the geometrical aspects of the field equation. Observationally, the data from Planck-2018 CMB [7, 8, 9, 10] proposes that $70\%$ of the energy budget in the universe is in the form of dark energy while the remaining $30\%$ of the energy budget of the universe is attributed to non-relativistic baryonic and non-baryonic pressureless ( $p_{m}=0$) dust matter with equation of state(EoS) $w_{m}=0$. In recent years, various candidates have been put forth to elucidate the nature of dark energy, with the cosmological constant standing out as one of the simplest models, embodying constant dark energy with an EoS of $w_{\Lambda}=-1$. Despite its alignment with observational data to a certain accuracy, the $\Lambda$CDM model grapples with unresolved cosmological constant and coincidence problems [16, 17, 18, 19, 20]. In response to these limitations, dynamical dark energy models, particularly those involving scalar fields, have gained prominence [21, 22, 23, 24, 25, 26] and with scalar field serving as a candidate of dynamical dark energy dubbed as $\phi$CDM having dynamical EoS $w_{\phi}$.
@@
In our paper, we have focused on the tachyonic field as a candidate for dark energy. A scalar field of this kind is driven by string theory as the negative-mass mode of the open string perturbative spectrum, though the primary use of it in the dark energy sector is phenomenological here in our paper, we have investigated the behaviour if tachyonic scalar field near infection point. Recently, the dark energy models with the inflection point in the context of inflation have been studied by several authors [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56].
To obtain a contemporary value of EoS $w_{\phi}\approx-1$ i.e. to get accelerated expansion mechanism of slow rolling is the key ingredient so, by imposing the approximation in which the kinetic term $\dot{\phi}^{2}/2$ of tachyon is significantly smaller than the potential $V(\phi)$ at the present moment i.e. $\dot{\phi}^{2}/2<<V(\phi)$ and this can be obtained by sitting $\phi$ in an extremely flat region of the potential.We consider the late time behavior of Tachyon scalar field in the accelerated expansion of the universe, where few similarities with early inflation. Numerous authors have examined the cosmic behaviour of the tachyonic scalar field [57, 58, 59], discovering possibilities for both pure exponential growth and power law expansion. We try to investigate the possibilities based on the superposition of these forms because the current data have not yet been able to determine the precise shape of the evolution of the scale factor.
@@ 1.1 Dynamical set up
$S==\frac{1}{2\kappa^{2}}\int d^{4}xf(\phi)\sqrt{-g}R+\int d^{4}xV(\phi)\sqrt{-g}\sqrt{\left(1+g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right)},$ (2)
@@ 2.1 For Case-I: $a(t)=\alpha t^{n}$
here $V_{0}=\frac{2n}{\kappa^{2}}\sqrt{1-\frac{2}{3n}}$ and $n>2/3$ for the expansion of the universe.
@@
From above Eq.(16) we found there is no such inflection point for this particular model and we can also see that from the Fig. (1) as well that the curve show no inflection but one can observe that the scalar field is rolling slowly along the potential. Now using Eq.(12) and on applying the slow roll approximation i.e. $\dot{\phi}\approx constant\implies\ddot{\phi}\approx 0$ we found the following evolution of scalar field with respect to time:
@@
here $t=t_{0}$ when $\phi=\phi_{0}$
Figure 2: The plot shows the evolution of $\phi$ with time $t$ for different values of the exponent factor $n$
Fig.(2) shows the evolution of $\phi$ for different values of $n.$
@@ 2.3 For Case-III: $a(t)=\eta t^{n}e^{\beta t}$
For this case we consider the universe in Quasi-exponential phase with the scale factor given by $a(t)=\eta t^{n}e^{\beta t}$ i.e. product of power law and exponential, here $\eta$, n and $\beta$ are the constants. For this quasi-exponential model the Hubble parameter is coming out to be:
@@
Now using Eq.(13), (14) and (20), one can found the following form of potential in terms of $t$ given as follow:
$V(t)=\frac{3}{\kappa^{2}}\sqrt{\left(\frac{n+\beta t}{t}\right)^{4}-\frac{2n}{3}\left(\frac{n+\beta t}{t^{2}}\right)^{2}}$ (21)
@@
$t=-\frac{n}{\beta}+\frac{n}{\beta}\exp{\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)}$ (22)
On substituting the Eq. (22) in (21) we got the following form of potential $V(\phi)$:
$V(\phi)=\frac{2n}{\kappa^{2}(\phi-\phi_{0})^{2}}\sqrt{\left(1+\sqrt{\frac{3}{2n}}\beta(\phi-\phi_{0})\right)^{2}-\frac{2}{3n}}$ (23)
On solving the above equation the inflection points for this particular potential are: $\phi_{0}=\frac{3\beta\phi\pm 2+6\sqrt{6n}}{3\beta}$
On substituting the value of potential (23) in equation of motion for this quasi-exponential tachyonic model, we get the following form of the evolution of scalar field:
$\phi(t)=\sqrt{2-2\left(\frac{\exp{(\frac{n+\beta t}{3\beta^{2}}+\frac{2-(\phi-\phi_{0})^{2}}{2})}}{(n+\beta t)^{n/3\beta^{3}}}\right)^{2}}+\phi_{0}$ (24)
Figure 3: The plot shows the evolution of $\phi$ with time $t$ for the first inflection point with different values of the exponent factor $n$ and $\beta$
Figure 4: The plot shows the evolution of $\phi$ with time $t$ for the second inflection point for the different values of the exponent factor $n$ and $\beta$
@@ 3 Dynamical Analysis:
Figure 5: phase portait of the dynamical system with $\lambda=1$ and $w=0$ near inflection point. Here point $A$ in the phase space shows the region where universe undergoes accelerated expansion
Figure 6: shows the phase portait for the above mentioned set of dynamical equations with $\lambda=1$ and $w=0$.
@@
Fig.(3) and (4) shows the phase portait of the derived set of autonomous equations and the region near the point A depict the accelerated expansion of the universe.
@@ 4 Conclusion
In this paper the evolution of minimally coupled Tachyonic scalar field at the inflection point has been investigated in an accelerating universe. The inflection point signifies a transition between regions of stability and instability. Near this point, the field’s dynamics has been investigated for different expansion factors. We have mainly considered the three categories of scale (expansion) factor namely: (a) Power law expansion ($a(t)=\alpha t^{n}$), (b) Exponential form ($a(t)=\gamma e^{\beta t}$) and (c) Quasi-Exponential expansion ($a(t)=\eta t^{n}e^{\beta t}$); for all these categories, the potentials and the corresponding possible inflection points have been calculated and the dynamical system of equations for the model has been obtained. Fig.(3) and (4) represents the critical points and the stability of the derived dynamical system of equations and the region near point A shows the accelerated expansion of the universe.

### v2

@@ Abstract
In this paper, we investigate the behaviour of a minimally coupled tachyonic scalar field at inflection point in an accelerating universe. We consider the different expansion factors and obtain potentials of tachyonic scalar field. Inflection points of homogeneous tachyonic scalar field is calculated for these potentials, then we employed the tools of dynamical system analysis for the considered potentials and obtained the stable points.
@@ 1 Introduction
The cosmic acceleration is now an established phenomenon. As supported by several cosmological and astrophysical observations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] our universe is in the phase of accelerated expansion at the present stage and generally, it is believed that behind this accelerated expansion there is a mysterious component with negative pressure dominating $70\%$ of the universe. One of the good candidates that can explain this acceleration is “Dark Energy” [11, 12, 13, 14, 15] which is introduced as modified matter in Einstein’s field equation which sits with the energy-momentum tensor. One of the remarkable features of dark energy is its negative pressure and repulsive gravity characteristics led by the equation of state $w_{de}<-1/3$. In alternative gravity theories, cosmic acceleration can be accounted for by modifications to the geometrical aspects of the field equation. Observationally, the data from Planck-2018 CMB [7, 8, 9, 10] proposes that $70\%$ of the energy budget in the universe is in the form of dark energy while the remaining $30\%$ of the energy budget of the universe is attributed to non-relativistic baryonic and non-baryonic pressureless ( $p_{m}=0$) dust matter with equation of state(EoS) $w_{m}=0$. In recent years, various candidates have been put forth to elucidate the nature of dark energy, with the cosmological constant standing out as one of the simplest models, embodying constant dark energy with an EoS of $w_{\Lambda}=-1$. Despite its alignment with observational data to a certain accuracy, the $\Lambda$CDM model grapples with unresolved cosmological constant and coincidence problems [16, 17, 18, 19, 20]. In response to these limitations, dynamical dark energy models, particularly those involving scalar fields, have gained prominence [21, 22, 23, 24, 25, 26], with scalar field serving as a candidate of dynamical dark energy dubbed as $\phi$CDM having dynamical EoS $w_{\phi}$.
@@
In our paper, we have focused on the tachyonic field as a candidate for dark energy. This type of scalar field is derived from string theory as the negative-mass mode of the open string perturbative spectrum. Though its primary use in the dark energy sector is phenomenological, we have investigated the behavior of the tachyonic scalar field near the inflection point. Recently, several authors have studied dark energy models with inflection points in the context of inflation. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57].
To obtain a contemporary value of EoS $w_{\phi}\approx-1$ which leads to accelerated expansion, the mechanism of slow rolling is a key ingredient. By imposing the approximation in which the kinetic term $\dot{\phi}^{2}/2$ of the tachyon is significantly smaller than unity i.e. $\dot{\phi}^{2}/2<<1$ this can be achieved by positioning $\phi$ in an extremely flat region of the potential. Potentials with an inflection point are good candidates for providing flat regions suitable for the slow roll approximation. We consider the late-time behavior of the tachyon scalar field in the accelerated expansion of the universe, which shares similarities with early inflation. Numerous authors have examined the cosmic behavior of the tachyonic scalar field [58, 59, 60], discovering possibilities for both pure exponential growth and power-law expansion. We try to investigate the possibilities based on the superposition of these forms because the current data have not yet determined the precise shape of the evolution of the scale factor.
@@ 1.1 Dynamical set up
$S=\frac{1}{2\kappa^{2}}\int d^{4}xf(\phi)\sqrt{-g}R+\int d^{4}xV(\phi)\sqrt{-g}\sqrt{\left(1+g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right)},$ (2)
@@ 2.1 For Case-I: $a(t)=\alpha t^{n}$
here $V_{0}=\frac{2n}{\kappa^{2}}\sqrt{1-\frac{2}{3n}}$.
@@
From above Eq.(16) we found there is no such inflection point for this particular model and we can also see that from the Fig. (1) as well that the curve show no inflection but one can observe that the scalar field is rolling slowly along the potential. Now using Eq.(12) and on applying the slow roll approximation i.e. $\dot{\phi}^{2}<<1\implies 1-\dot{\phi}^{2}\approx 1$ we found the following evolution of scalar field with respect to time:
@@
here $t=t_{0}$ when $\phi=\phi_{0}$.
Figure 2: The plot shows the evolution of $\phi$ with cosmological redshift $(1+z)$ for different values of the exponent factor $n$; with $\alpha=0.1$ and $\phi_{0}<<1$.
In terms of cosmological redshift we can also rewrite the above Eq.(17) as:
$\phi(1+z)=\sqrt{\frac{2}{3n\alpha^{2/n}}}\sqrt{\left(\left(\frac{1}{1+z}\right)^{2/n}-\left(\frac{1}{1+z_{0}}\right)^{2/n}\right)}+\phi_{0}$ (18)
Fig.(2) shows the evolution of $\phi$ for different values of $n$. We examine the behavior of a tachyonic scalar field, $\phi$, as the redshift, $(1+z)$, approaches zero, $\phi(1+z)$ approaches infinity corresponding to the late-time evolution of the universe. The tachyonic field is characterized by a unique potential which is derived from the scale factor, this potential governs its dynamics and the corresponding impact on cosmic expansion. As the universe evolves towards the present era, the tachyonic field enters a phase where its kinetic energy becomes negligible, and it slowly rolls down its potential. This slow-roll regime results in the field behaving similarly to a cosmological constant, with a near-constant energy density. As a consequence, the universe continues to expand at an accelerating pace.
@@ 2.3 For Case-III: $a(t)=\eta t^{n}e^{\beta t}$
For this case we consider the universe the mixed form of the scale factor given by $a(t)=\eta t^{n}e^{\beta t}$ i.e. product of power law and exponential terms, here $\eta$, $n$ and $\beta$ are the constants determining overall rate at which the scale factor changes. For this model the Hubble parameter is obtained as:
@@
Now using Eq.(13), (14) and (21), one can found the following form of potential and the scalar field in terms of $t$ as follow:
$V(t)=\frac{3(n+\beta t)^{2}}{\kappa^{2}t^{2}}\sqrt{1-\frac{2n}{3(n+\beta t)^{2}}}$ (22)
@@
$t=-\frac{n}{\beta}+\frac{1}{\beta}\exp{\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)}$ (23)
On substituting the Eq. (23) in (22) we got the following form of potential $V(\phi)$ which depends in the scalar field $\phi$ as:
$V(\phi)=\frac{3\beta^{2}\exp{\left(2\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)}}{\left(\exp{(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0}))}-n\right)^{2}}\sqrt{1-\frac{2n}{3\exp{\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)^{2}}}}$ (24)
By using binomial expansion and on simplification, Eq.(24) can be rewritten as:
$V(\phi)=3\beta^{2}\left(\frac{\exp\left(2\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)-n/3}{\left(\exp{\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)}-n\right)^{2}}\right)$ (25)
here for the simple case we use the following assumption:
$\exp{\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)}\approx 1+\left(\beta\sqrt{\frac{3}{2n}}(\phi-\phi_{0})\right)$ (26)
and solving above Eq.(25) using the approximation mentioned above, one can easily find that the potential shows inflection at:
$\phi=\frac{2}{3}\sqrt{\frac{2n}{3\beta^{2}}}\left(\frac{(n-4)(2n-3)}{(n-3)(n-8)}\right)+\phi_{0}$ (27)
For $n=3$ or $n=8$ inflection point approaches infinity and for the case when $n=4$ or $n=3/2$ inflection point becomes $\phi\longrightarrow\phi_{0}$. On substituting the value of potential (24) in equation of motion for this tachyonic field model, for the simplest case of $n=1$, we get the following form of the evolution of scalar field with respect to cosmological redshift at inflection point:
$\phi(1+z)=\frac{4}{\beta}\sqrt{\frac{3}{2}}\left(\frac{1}{4\eta(1+z)}-\log{\left(\frac{1}{4\eta(1+z)}\right)}\right)$ (28)
Fig.(3) shows the evolution of scalar field with respect to the redshift $(1+z)$. From the Fig.(3) and Eq.(28) we can see that the scalar field evolves asymptotically as $(1+z)\longrightarrow 0$ signifying the phase of accelerated expansion. Using Eq.(28) and Eq.(11), the EoS (Equation of State) parameter of this inflection point tachyonic scalar field model $w_{\phi}\approx(0.4(1-4\eta(1+z))-1)\longrightarrow-0.8$ as $(1+z)\longrightarrow 0$ with $\beta\approx 0.05$ supporting eternal accelerated expansion of the universe at late time.
Figure 3: The plot shows the evolution of $\phi$ with cosmological redshift $(z+1)$ at inflection point with different values of the exponent factor $\beta$ for the simplest case when $n=1$ and $\eta=1$.
@@ 3 Dynamical Analysis:
Figure 4: phase portait of the dynamical system with $\lambda\longrightarrow 1$ and $w=0$ near inflection point. Here point $A$ in the phase space shows the region where universe undergoes accelerated expansion
@@
Fig.(4) shows the phase portait of the derived set of autonomous equations and the region near the point A depict the accelerated expansion of the universe.
@@ 4 Conclusion
In this paper the evolution of minimally coupled Tachyonic scalar field at the inflection point has been investigated in an accelerating universe. The inflection point signifies a transition between regions of stability and instability and inflection point provides a flat region suitable for the slow roll approximation. Near this point, the field’s dynamics has been investigated for different forms of expansion factors. We have mainly considered the three categories of scale (expansion) factor namely: (a) Power law expansion ($a(t)=\alpha t^{n}$), (b) Exponential form ($a(t)=\gamma e^{\beta t}$) and (c) Mixed form ($a(t)=\eta t^{n}e^{\beta t}$); for all these categories, the potentials and the corresponding possible inflection points have been calculated and the dynamical system of equations for the model has been obtained. Out of these three categories, mixed form shows inflection at $\phi$ given in Eq.(27) and near this point under slow roll approximation the evolution of scalar field with cosmological redshift $(1+z)$ shows the accelerated expansion of the universe at late times. In section (3) dynamical analysis carried out shows at critical point $A(1/5,\sqrt{3}/5\lambda)$ as shown in Fig.(4) and (LABEL:fig:4b) represents the stability of the derived dynamical system of equations and the region near inflection point which shows the accelerated expansion of the universe.